OpenStudy (anonymous):

Find the value of the expression 2x3 + 3y2 − 17 when x = 3 and y = 4.

Vocaloid (vocaloid):

x = 3, y = 4 so we substitute these values into the equation to get 2x^3 + 3y^2 - 17 = 2(3)^3+3(4)^2 - 17 = ?

OpenStudy (anonymous):

thank you sooo much